### v2

# Ramanujan’s function on small primes

Abstract.
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number. For this reason, as an example in a recent article about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime. We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit regularities of the following kind: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in a nearly periodic fashion with $n$.
Key words and phrases:
partitions, graphs, cusp forms
2020 Mathematics Subject Classification:
05A17, 05C62, 11F11, 11P83
11 December 2025
1. Introduction
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47] . For this reason, we made computer calculations about the function on positive integers
$\tau_{p}:n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime. Using matrix machinery reproduced in Lemmas 2.1 and 2.2 below, within the range of our observations, the eigenvalues of the matrices $H_{n}$ and $J_{n}$ attached to this function by the two lemmas behave as follows: the minimum modulus among the eigenvalues for each of the two matrices associated to a particular $\tau_{p}(n)$ appear to oscillate with $n$. The connection to Lehmer’s question is that the associated matrix $H_{n}$ has a zero eigenvalue if and only the associated matrix $J_{n}$ does too, and, if they do, $\tau_{p}(n)=0$.
We show plots of the oscillations in Figure 1 and Figure 3 below. We also describe some pleasant patterns among the abscissas of the points on their lower envelopes that seem unlikely to persist unaltered: they seem to suggest that the effect of Ramanujan’s function on the prime sequence is to tune somehow the noise in it, and we do not know how that might work. We have looked at similar compositions of tau with other functions and, for them, this “tuning” is not so clear-cut.
2. Lemmas
In the context of the theory of symmetric functions, the lemmas in this section are well known. They appear for example, in MacDonald’s book [M]), but the proofs are not easy to locate. (In fact, we have not located them.) We have deposited home-made proofs of these statements in our depository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”. In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
Lemma 2.1.
The equations below are equivalent. (We will refer to both of them as equation (D) in the sequel.) Let $h_{0}=1$ and
$nh_{n}=\sum_{r=1}^{n}j_{r}h_{n-r}$
for $n\geq 1$. With $H(t)=\sum_{n=0}^{\infty}h_{n}t^{n}$ and $J(t)=\sum_{r=1}^{\infty}j_{r}t^{r}$,
$t\frac{d}{dt}H(t)=H(t)J(t).$
With $f_{0}$ possibly defined as equal to one,, let $\overline{f}$ denote the sequence $\{f_{n}\}_{n\in\mathbb{Z}^{+}}$. Let $J_{n}(\overline{j})$ and $H_{n}(\overline{h})$ be the matrices
$\left(\begin{matrix}j_{1}&-1&0&\cdots&0\\ j_{2}&j_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ j_{n-1}&j_{n-2}&\cdots&j_{1}&-n+1\\ j_{n}&j_{n-1}&j_{n-2}&\cdots&j_{1}\\ \end{matrix}\right)$
and
$\left(\begin{matrix}h_{1}&1&0&\cdots&0\\ 2h_{2}&h_{1}&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ nh_{n}&h_{n-1}&h_{n-2}&\cdots&h_{1}\\ \end{matrix}\right),$
respectively.
Lemma 2.2.
Let $h_{0}=j_{1}=1$. Equation (D) implies that
(1) $j_{n}=(-1)^{n+1}|H_{n}(\overline{h})|.$
and
(2) $n!h_{n}=|J_{n}(\overline{j})|$
for $n\geq 1$.
Next we state a corollary of the following theorem in Vein and Dale’s book  [Vein1999].
Proposition 1.
(Vein and Dale, Theorem 4.23) Let $A_{n}=$
$\begin{pmatrix}a_{1}&-1&0&\cdots&0\\ a_{2}&a_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n-1}&a_{n-2}&\cdots&a_{1}&-(n{-}1)\\ a_{n}&a_{n-1}&\cdots&a_{2}&a_{1}\end{pmatrix}^{T},$
where the $T$ operator is matrix transpose. Let $B_{n}(x)=|A_{n}-xI_{n}|$, so that $B_{n}(x)=(-1)^{n}\chi_{A_{n}}(x)$. Then
$B_{n}(x)=\sum_{r=0}^{n}\binom{n}{r}|A_{r}|x^{n-r}.$
Theorem 2.3.
 [^2] Here we are merely codifying an argument shown to us by J. López-Bonilla.
With $h$ and $j$ as in the above lemmas,
$\chi({(J_{n}(\overline{j}))}(x)=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{r}x^{n-r}.$
Proof.
Determinants are invariant under the transpose, so, setting each $a_{i}$ equal to $j_{i}$ in the lemmas above, the proposition can be restated in the following way: Let $J_{n}(\overline{j})=$
$\begin{pmatrix}j_{1}&-1&0&\cdots&0\\ j_{2}&j_{1}&-2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ j_{n-1}&a_{n-2}&\cdots&j_{1}&-(n{-}1)\\ j_{n}&j_{n-1}&\cdots&j_{2}&j_{1}\end{pmatrix}$.
Let $C_{n}(x)=|(J_{n}(\overline{j})-xI_{n}|$, so that $C_{n}(x)=(-1)^{n}\chi({(J_{n}(\overline{j}))}(x)$. By Proposition 1, $C_{n}(x)=$
$\sum_{r=0}^{n}\binom{n}{r}|(J_{r}(\overline{j})|(-1)^{n-r}x^{n-r}=$
(by Lemma 2.2)
$\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-x)^{n-r}=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{n-r}x^{n-r}.$
Simplifying, $\chi({(J_{n}(\overline{j}))}(x)=\sum_{r=0}^{n}\binom{n}{r}r!h_{r}(-1)^{r}x^{n-r}.$
3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
3.1. Characteristic functions of the undeformed matrices of the form $J_{n}(\overline{j_{\tau_{p}}})$ representing the function $\tau_{p}$.
Let $h_{\tau_{p}}(0)=1$ and $h_{\tau_{p}}(n)=\tau_{p}(n)$ for positive integers $n$. Let $j_{\tau_{p}}(n)$ and $h_{\tau_{p}}(n)$ together satisfy equation (D). Given $h_{\tau_{p}}$, the elements of $\overline{j_{\tau_{p}}}=\{j_{\tau_{p}(n)}\}_{n=1,2,...}$ can be computed from equation (D1). By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j_{\tau_{p}}})|/n!$. Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau_{p}}}))(x)=|xI-J_{n}(\overline{j_{\tau_{p}}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau_{p}}})$. The following statements are clearly equivalent: (a) $\tau(p_{n})=0$,
(b) $|J_{n}(\overline{j_{\tau_{p}}})|=0$, and (c) $\Pi_{n}(0)=0$.
Writing [^3] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25]. $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k}$, Theorem 2.3 gives
(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
In (theoretical) practice, this result would allow us to bypass the construction of the matrices $J_{n}(\overline{j_{\tau_{p}}})$ in our computations, saving time and allowing us to extend our data sets. We have not found occasion to exploit this advantage; if a similar result applied to the $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))(x)$, we would have used it; at present, this is not available to us.
Remark 1.
Let $M$ be an $n\times n$ matrix over a field. For any subset $S\subseteq\{1,2,\ldots,n\}$, let $M[S,S]$ denote the submatrix of $M$ with rows and columns indexed by $S$. The following proposition is well known. (For example, see equation (1.2.13) in Horn and Johnson’s book  [Horn2012].)
Proposition 2.
The coefficient $a_{k}$ of $x^{k}$ in $\chi(M)$ satisfies
$a_{n-k}=(-1)^{k}\sum_{\begin{subarray}{c}\#S=k\\ {S\subseteq\{1,2,\ldots,n\}}\end{subarray}}|M[S,S]|.$
The number of terms in the sum is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$. But this turns out to be too good to be true. [^4] See ‘too good.ipynb’in [Bre25]. Instead, there must be systematic cancellations coming from those symmetries. So far, we have not understood these cancellations.
3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau_{p}}})$.
[…]
Let $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))=\Pi_{n}^{(1)}(x)$ (say.) (The corresponding objects for $c=0$ exhibit behaviors similar to those we go over here. [^5] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].) Let $V_{n}^{(1)}$ be the set of zeros of $\Pi_{n}^{(1)}(x)$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ lies on an oscillating curve.
Figure 1. Minimum moduli of the roots of the $\Pi^{(1)}_{n}(x)$
and their logarithms.
Figure 2. Minimum moduli of the roots of undeformed $\Pi_{n}(x)$.
The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the roots of $\Pi_{n}^{(1)}(x)$ to the origin may behave in a predictable way, but the details are not clear to us because (among other reasons) our data is limited. Also, we do not know the details of the relationship of the roots of $\Pi_{n}(x)$ to those of the $\Pi_{n}^{(c)}(x)$, so the connection, if any, to Lehmer’s question is also unclear. Furthermore, we have no statement like equation (3) for the $\Pi_{n}^{(c)}(x)$, because we do not have an explicit description of the sequence $\{h_{n}\}$ represented in Lemma 2.2 by the matrices of which the $\Pi_{n}^{(c)}(x)$ are the characteristic polynomials. As we said above, such a description would permit us to sidestep the construction of the matrices we currently use to find these polynomials and the resulting time savings would in practice allow us to collect more data. More data would be welcome in view of the oscillations we are trying to check.
The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 1 appear to behave regularly: the abscissas of the points on the lower envelope form a set $\{5,9,...,395\}$, which, when they are replaced by their residues modulo $4$, make Table 1.
1 1 1 1 1 1 1 1 1
1 2 2 2 2 2 2 2 2
2 3 3 3 3 3 3 3 3
3 0 0 0 0 0 0 0 0
0 1 1 1 1 1 1 1 1
1 2 2 2 2 2 2 2 2
2 3 3 3 3 3 3 3 3
3 0 0 0 0 0 0 0 0
0 1 1 1 1 1 1 1 1
1 2 2 2 2 2 2 2 2
2 3 3 3 3 3
Table 1. Residues modulo 4 of abscissas from the lower envelope of Figure 1.
The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others. [^6] We are using data from the lower envelope of the logarithms of the minimum moduli. (The lower plot in Figure 1.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25].
For contrast, we reproduce in Figure 2 a plot of the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$. We have not yet attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$. The irregularities visible in Figure 2 do not seem promising to us in that regard. This is an obstacle to using these plots to look at Lehmer’s question because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the function $\tau_{p}$ in a straightforward way.
3.3. Characteristic functions of deformed matrices $H_{n}^{(1)}(\overline{h_{\tau_{p}}})$.
 [^7] See our notebook “tauprime=j H matrix4dec25no2.ipynb” in [Bre25].
Given the sequence $\{\tau_{p}(n)\}_{n=1,2,...}=$ (say) $\{j_{\tau_{p}}(n)\}_{n=1,2,...}=\overline{j_{\tau_{p}}}$, we
construct a sequence $\overline{h_{\tau_{p}}}=\{h_{\tau_{p}}(n)\}_{n=1,2,...}$ by solving the convolution (D1). By Lemma 2.2, $\tau_{p}(n)=(-1)^{n+1}|H_{n}(\overline{h})|$. Then we construct matrices
[…]
Figure 3. Minimum moduli and their logarithms for zeros of the $\Psi^{(1)}_{n}(x)$.
The table of modulo $4$ residues for the abscissas of the lower envelopes of the plots in Figure 3 is not as regular as the corresponding table for Figure 1; the lengths of the strings of constant residue vary:
0 0 2 2 3 3 3 3 3
0 0 0 0 0 1 1 1 1
1 1 2 2 2 2 2 3 3
3 3 3 3 0 0 0 0 0
1 1 1 1 1 1 2 2 2
2 2 2 2 3 3 3
Table 2. Residues modulo 4 of abscissas from thelower envelope of Figure 3.